0.25=(1-r^2)^2+(0.73r)^2, what is r equal to?
Answer #1

2 i 8=? 0.25 (1-x2) +(00738) 0.25 = 1+ (v2) 2 - 2 8² + 6.7322 let 82=2 - x ² - 10467/x+0.75 = 0 = x= 0.73355+ 0.46033 i =) 8